# Multilevel Markov Chain Monte Carlo with likelihood scaling

### v2

@@ 6 Conclusion
We show how the existing multilevel Markov chain Monte Carlo methodology, which has previously been used effectively for high-resolution models with nevertheless low-resolution discrete observations, can be extended to fit in the case where both the model and the data are of very high resolution. We do this by extending the level-dependent treatment of parameters and discretisations to also include the observations. We add a weighting function to the observations which reduces their dimensionality such that these weighted observations can be used as proxies to the true observations at coarser levels in the multilevel algorithm.

### v3

@@ 6 Conclusion
We show how the existing multilevel MCMC methodology, which has previously been used effectively for high-resolution models with nevertheless low-resolution discrete observations, can be extended to fit in the case where both the model and the data are of very high resolution. We do this by extending the level-dependent treatment of parameters and discretisations to also include the observations. We add a weighting function to the observations which reduces their dimensionality such that these weighted observations can be used as proxies to the true observations at coarser levels in the multilevel algorithm.